OpenStudy (anonymous):

check my work Solve for x: 4 over x plus 4 over quantity x squared minus 9 equals 3 over quantity x minus 3 am I right? x = 4 and x = -9

OpenStudy (anonymous):

is it suppose to be \[(\frac{ 4 }{ x }) + 4x^2 -9 = \frac{ 3 }{ x-3 }\]

OpenStudy (raden):

actually, you can write it using Latex @ninashabrang

OpenStudy (anonymous):

can solve me it though it so I can see it..?

OpenStudy (anonymous):

well we are confused witht the question that you wrote can ou just write down on draw or latex

OpenStudy (anonymous):

ill draw it hold on

OpenStudy (anonymous):